Suppose a small economy has consumption spending of $6,500 when its income is $10,000. When income goes up to $11,000 then consumption spending is $7,250. What is the government spending multiplier in this economy? (Hint: Find the MPC first, then calculate the multiplier.) 2.85 4.00 1.53 7.00
Answer #1

Answer : 4.00

There are three methods, we can calculate the national income (Y) - a) product method,b) income method, and c) expenditure method.

The expenditure method adds the total expenditure in an economy to of calculate the national income. It is the sum total of consumption expenditure(C) by household sector on goods and services,investment expenditure (I), government expenditure (G) on goods and services, and net exports (X-M) in an economy.

Y = C + I + G + (X-M)

The increase in any one of the variables above increases the national income. When the government increases its spending or expenditure (G), it will have an expansionary effect on national income (Y).As a result the national income rises. The change in national income due to the change in government spending is called the government spending multiplier.It is denoted by 'Y / G'.

Y / G = 1/ 1- MPC , where 'MPC' is the marginal propensity to consume , '' stands for change.

MPC is the increase in consumption spending for the increase in income or more precisely for the increase in disposable income (income after tax payment).

MPC = C / Y

In the question above,

when Y = $10,000 , C = $6,500

when, Y = $11,000 , C = $7,250

Y = $11,000 - $10,000 = $1000

C = $7,250 - $6,500 = $750

Now, MPC = C / Y = $750 / $1000

Or, MPC = 0.75

1 - MPC = 1 - 0.75 = 0.25

Y / G = 1/ (1- MPC ) = 1/0.25 = 4.00

So, the government spending multiplier in this economy is 4.00 , i.e., if the government spending rises by $100, its expansionary effect on national income will be four times. So income will rise by (4 x 100 = $400)

______________________________________________________________________